

********************************************************************
UARS WEEKLY STATUS REPORT - 20 Feb 2004
********************************************************************
***********
SOLSTICE
***********
During this week SOLSTICE continued to operate in good health gathering
solar data on all available orbital tracking periods. On average, we
received 53% of each day's data amounting to about 350 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered all of its daily mid (1.1 nm)
resolution, Mg II, occultation, solar pointing offset and high
resolution continuous monitoring scans during this past week for
periods of UARS real-time telemetry or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003149 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1993299. MEPS data files between
1993300 and 1993329 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 07 February 2004 (Orbit 67886, GMT Day 038) through
13 February 2004 (Orbit 67990, GMT Day 044).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
NONE THIS REPORT PERIOD
The beta angle ranged between the angles of 40.8 degrees to 27.3
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 21 February 2004 (DOY 052).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.27 to 1.29 degrees C. Additional
battery performance data for each battery on 06 February and 13
February 2004 are provided in Appendix D.
The clock error ranged from 10.1 to -11.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The problem with the new CMS code was found this week. Two local
variables, which effectively set the length of the ephemeris tables,
were deleted and put into a global variable. This change caused both
tables to be built to the length of the TDRS ephemeris. Thus the
spacecraft ephemeris load was incorrectly sized and produced the
signature of good data then zero fill, which was seen in the anomalous
load(s). These variables were fixed and a new release of the CMS code
was delivered to the CMS backup machine. While this version is being
tested, the FOT will build all loads on the old version of the CMS
code, except the 3-TDRS ephemeris that must be built on the new
version. The FOT will also continue to perform their manual checks to
insure the ephemeris loads are being built correctly.
MEETINGS
The FOT held a meeting on 09 February 2004 to discuss the CMS code
problem and the results of the analysis performed the previous week.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 February 2004 (DOY 052).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 13 February 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
67918 040/0330 04-052 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67936 041/0729 04-053 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 13 February 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 13 February 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
67918 040/0332 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67936 041/0734 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 13 February 2004
06 February 2004 - GMT Day 037 Beta = 40.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 +16.8/-145.6 +5.47/+4.51 +26.8/-15.2 25.1 1.03 20.2
3 +11.2/-100.8 +4.51/+3.23 +26.4/-15.2 25.1 1.02 19.9
13 February 2004 - GMT Day 044 Beta = 27.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-3.99 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-089.6 +6.43/+5.14 +28.8/-15.2 24.0 1.04 20.7
3 +11.2/-134.4 +5.14/+3.87 +28.0/-15.2 24.0 1.03 20.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 13 February 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-010 FOT 67934 041/0409 35016 4 Sec Data Dropout
04-011 FOT 67942 041/1747 30832 Generic Late Acquisition
04-012 FOT 67965 043/0643 35015 Late Acq & No GCMRs/UPDs
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-001 FOT See List Jan 04 30832 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-009 FOT 67854 035/2126 N/A EPHEMERIS REJECTED
04-011 FOT See List Feb 04 30832 Generic Late Acquisition
04-010 FOT 67934 041/0409 35016 4 Sec Data Dropout
04-012 FOT 67965 043/0643 35015 Late Acq & No GCMRs/UPDs